# Solar Cruiser Disturbance Torque Estimation and Predictive Momentum Management

### v1

@@ Abstract
This paper presents a novel disturbance-torque-estimation-augmented model predictive control (MPC) framework to perform momentum management on NASA’s Solar Cruiser solar sail mission. Solar Cruiser represents a critical step in the advancement of large-scale solar sail technology and includes the innovative use of an active mass translator (AMT) and reflectivity control devices (RCDs) as momentum management actuators. The coupled nature of these actuators has proven challenging in the development of a robust momentum management controller. Recent literature has explored the use of MPC for solar sail momentum management with promising results, although exact knowledge of the disturbance torques acting on the solar sail was required. This paper amends this issue through the use of a Kalman filter to provide real-time estimation of unmodeled disturbance torques. Furthermore, the dynamic model used in this paper incorporates key fidelity enhancements compared to prior work, including the Solar Cruiser’s four-reaction-wheel assembly and the offset between its center of mass and center of pressure. Simulation results demonstrate that the proposed policy successfully manages angular momentum growth under slew maneuvers that exceed the operational envelope of the current state-of-the-art method. The inclusion of the disturbance torque estimate is shown to greatly improve the reliability and performance of the proposed MPC approach. This work establishes a new benchmark for Solar Cruiser’s momentum management capabilities and paves the way for MPC-based momentum management of other solar sails making use of an AMT and/or RCDs.
@@ 1 Introduction
Prior work by Shen and Caverly (2025b) developed MPC-based momentum management strategies tailored for solar sails equipped with an AMT and RCDs. They developed a dynamic model that captured key features of the AMT movement, including the resultant time-varying changes in the spacecraft’s CM and moment of inertia matrix. The MPC policy developed in the work of Shen and Caverly (2025b) leveraged its optimization capabilities to handle the actuator constraints and requirements, including the enforcement of on-off RCD actuation and AMT motion rate limits, all while incorporating tuning parameters designed to adjust the trade-off between system performance and control effort. For real-time onboard implementation, the MPC formulation was posed as a quadratic program (QP), which guarantees fast and robust convergence suitable for the short processing cycles required by the flight computer. The work of Shen and Caverly (2025a) further examined the critical balance between model fidelity and computational cost when implementing the MPC policy developed by Shen and Caverly (2025b) to determine feasibility for real-time implementation. However, these MPC implementations both make an unrealistic assumption that exact knowledge of the disturbance torque acting on the solar sail is available for use within MPC’s prediction model. This is a significant assumption that limits practical implementation of the MPC policy proposed by Shen and Caverly (2025b), as unpredictable shape deformation of the solar sail and temporal changes in the sail’s optical properties make it virtually impossible to accurately predict disturbance torques from analytical models (Wang et al., 2025; Gauvain and Tyler, 2023). The use of an inaccurate torque model within the MPC framework significantly degrades momentum management performance, negating the purported benefits of the MPC momentum management policy. Another practical limitation of the work of Shen and Caverly (2025b, a) is that their implementations assume that the solar sail is equipped with three RWs aligned with the principal axes of its body-fixed frame. Many spacecraft, including Solar Cruiser, have a 4-RW assembly for redundancy and increased performance, which precludes the use of the methods developed by Shen and Caverly (2025b, a). Furthermore, the implementation in the work of Shen and Caverly (2025b) assumed that the RWs remained in the same plane as the solar sail’s CP, which is not representative of Solar Cruiser’s geometry.
@@
To overcome the limitations of prior work, this paper presents a novel MPC-based momentum management policy that incorporates disturbance torque estimation and a 4-RW assembly tailored for the Solar Cruiser mission. A Kalman filter framework is used to estimate the unmodeled disturbance torques and system model errors in real time, thus enhancing the predictive capability of the MPC. Similar Kalman-filtering approaches have been used in the literature to estimate unknown parameters or terms within a system model (Zenere and Zorzi, 2018; Woodbury and Junkins, 2010; Hayes and Caverly, 2025; Ahmed et al., 2024). For example, Hayes and Caverly (2025) used a Kalman filter to estimate the unknown atmospheric density of a satellite during an orbital reentry, while Ahmed et al. (2024) estimated the unknown wind acting on a small uncrewed air vehicle. Solar Cruiser’s 4-RW assembly is accounted for within the proposed MPC implementation through the use of the commonly-used pseudo-inverse RW allocation approach (Leve et al., 2015; Markley and Crassidis, 2014). This provides the MPC prediction model with accurate knowledge of the dynamics of each individual RW, allowing for their operation to be constrained within their saturation limits.
@@
This paper presents three key contributions relative to the state-of-the-art in solar sail momentum management, including prior work on MPC-based methods by Shen and Caverly (2025b, a). The first contribution is a robust MPC momentum management formulation that uses a Kalman filter to estimate unknown disturbance torques actin on the solar sail. To the best of the knowledge of the authors, this is the first realistically-implementable momentum management policy for a solar sail equipped with an AMT and RCD that outperforms the method of Tyler et al. (2023). The second contribution is the incorporation of a 4-RW assembly within an MPC-based momentum management policy. To the best of the knowledge of the authors, this is the first MPC-based momentum management policy to consider a realistic 4-RW assembly. The third contribution is an assessment of the proposed momentum management policy in a realistic simulation of Solar Cruiser’s dynamics, where its CM is located a distance from the sail plane and non-ideal SRP forces are considered. Both these effects are meaningful when considering Solar Cruiser’s dynamics, as they result in substantial coupling between axes, yet they were not considered in the prior work of Shen and Caverly (2025b).
@@
Details of the nonlinear system dynamics of Solar Cruiser, the 4-RW control allocation algorithm, and the momentum management actuators are presented in Section 2. This section also provides the linearized dynamic model used in the Kalman filter and MPC frameworks. The Kalman filter formulation is presented in Section 3, providing details of how the unmeasurable disturbance torques are estimated. The MPC formulation is presented in Section 4, detailing the implementation of estimation-prediction framework and the incorporation of the 4-RW assembly into the MPC prediction model. Numerical simulation results are presented in Section 5, validating the performance of the proposed estimation-augmented MPC with comparisons to Solar Cruiser’s state-of-the-art momentum management method (Tyler et al., 2023). Results in this section are also presented that demonstrate the effect of actuation thresholds within the proposed momentum management policy on actuation efficiency and observability of the roll-axis disturbance torque.
@@ 2 Attitude Dynamics and Control Actuation
NASA’s Solar Cruiser uses AMT and RCDs as its momentum management actuators (Inness et al., 2023). The AMT changes the relative alignment of the CM and CP such that the SRP force acting on the CP results in a corresponding torque about the CM. Moving the AMT and appropriately placing the CM/CP offset results in a controllable moment that unloads the accumulated RW angular momentum. However, the moment of inertia matrix changes when the AMT moves and the mass distribution of the sailcraft changes. The dynamics are thus coupled with the AMT translation. This section presents the dynamics and control of the sailcraft, starting with important notation and proceeding with its attitude dynamics, the RW attitude control law, and details regarding the momentum management control actuation.
@@ 2.1 Notation
The direction cosine matrix (DCM) $\mbf{C}_{ba}$ describes the attitude of reference frame $\mathcal{F}_{b}$ relative to reference frame $\mathcal{F}_{b}$. While different attitude parameterizations can be used to describe a DCM, Euler-angle sequences are used in this work due to their ease of physical interpretation and the lack of any large-angle maneuvers that would potentially result in a kinematic singularity. The DCM can be used to express a physical vector in different reference frames. For example, $\mbf{v}_{b}=\mbf{C}_{ba}\mbf{v}_{a}$.
@@ 2.2 Solar Cruiser Attitude Dynamics
Following the approach of Shen and Caverly (2025b) and as illustrated in Fig. 1, Solar Cruiser’s sail is modeled as a thin flat plate (denoted as $\mathcal{S}$) and a rigid rectangular bus (denoted as $\mathcal{P}$). The nonlinear rigid-body attitude dynamics of a solar sail incorporating the AMT translation as a control input were derived by Shen and Caverly (2025b), including the time-dependent moment of inertia matrix corresponding to the AMT position. It is assumed that the RWs are colocated at the CM of the bus (point $p$) since the dimension of the sail $\mathcal{S}$ is much greater than the dimension of the bus $\mathcal{P}$. As shown in Fig. 1, the sailcraft body frame $\mathcal{F}_{b}$ is defined with $\underrightarrow{b}^{3}$ pointing through the normal (roll) axis of the sail, and $\underrightarrow{b}^{2}$, $\underrightarrow{b}^{1}$ pointing within the plane of the sail and representing the pitch and yaw axes, respectively. The moment of inertia matrix relative to the sailcraft’s CM (point $c$) is defined as
@@
Let $\mathcal{F}_{a}$, defined by basis vectors $\underrightarrow{a}^{1}$, $\underrightarrow{a}^{2}$, $\underrightarrow{a}^{3}$, be an inertial reference frame. The solar sail attitude dynamics with mass translation are defined as (Shen and Caverly, 2025b)
@@
where ${\boldsymbol{\omega}}_{b}^{ba}$ is the angular velocity of $\mathcal{F}_{b}$ relative to $\mathcal{F}_{a}$ resolved in $\mathcal{F}_{b}$, $\mbf{h}_{b}^{\text{RW}}$ is the collective angular momentum of the RWs relative to the CM of the bus (point $p$) with respect to inertial reference frame $\mathcal{F}_{a}$ resolved in $\mathcal{F}_{b}$. The term ${\boldsymbol{\tau}}_{b}^{\mathcal{B}c}={\boldsymbol{\tau}}_{b}^{\text{AMT}}+{\boldsymbol{\tau}}_{b}^{\text{RCD}}+{\boldsymbol{\tau}}_{b}^{\text{dist}}$ is the torque acting on the sailcraft relative to its CM, consisting of the AMT-SRP-induced torque ${\boldsymbol{\tau}}_{b}^{\text{AMT}}=\frac{m_{s}}{m_{p}+m_{s}}\mbf{r}_{b}^{ps^{\times}}\mbf{f}_{b}^{\text{SRP}}$, the RCD torque ${\boldsymbol{\tau}}_{b}^{\text{RCD}}$, and the disturbance torque ${\boldsymbol{\tau}}_{b}^{\text{dist}}$. The term $\mbf{f}_{b}^{\text{SRP}}$ denotes the solar radiation pressure induced force acting on the sail’s CP. It is assumed that SRP force $\mbf{f}_{b}^{\text{SRP}}$ and disturbance torque ${\boldsymbol{\tau}}_{b}^{\text{dist}}$ are constant, which is justified given the slow evolution of the sailcraft attitude in this work. The RCD torque is generated by activating one of the set of four RCDs (either brown or orange) visualized in Fig. 1. Actuating the orange RCDs generates a positive roll torque about $\underrightarrow{b}^{3}$ axis, while actuating the brown RCDs generates a negative torque about this same axis. The time-dependencies of $\mbf{J}_{b}^{\mathcal{B}c}(t)$, $\mbf{r}^{ps}_{b}(t)$, ${\boldsymbol{\omega}}_{b}^{ba}(t)$, $\mbf{h}_{b}^{\text{RW}}(t)$, ${\boldsymbol{\tau}}_{b}^{\mathcal{B}c}(t)$ are omitted for brevity in Eq. (1).
@@
A 3-2-1 Euler angle sequence is used to describe the rotation between $\mathcal{F}_{a}$ and $\mathcal{F}_{b}$, so that $\mbf{C}_{ba}=\mbf{C}_{1}(\theta_{1})\mbf{C}_{2}(\theta_{2})\mbf{C}_{3}(\theta_{3})$ is the DCM describing the orientation of $\mathcal{F}_{b}$ relative to $\mathcal{F}_{a}$, where $\mbf{C}_{i}(\cdot)$ is the DCM representing a rotation about the $i$-th principal axis. The mapping ${\boldsymbol{\omega}}^{ba}_{b}=\mbf{S}({\boldsymbol{\theta}})\dot{{\boldsymbol{\theta}}}$ relates Euler angle rates to angular velocity, where the matrix ${\boldsymbol{\theta}}=\Big[\theta_{1}\,\,\,\theta_{2}\,\,\,\theta_{3}\Big]^{\mathsf{T}}$ is the set of Euler angles and $\mbf{S}({\boldsymbol{\theta}})$ is the mapping matrix. It is worth noting that $\mbf{S}({\boldsymbol{\theta}})$ depends only on $\theta_{1}$ and $\theta_{2}$ due to the selected 3-2-1 Euler angle sequence. Given that a solar sail is designed to keep the Sun within its field of view and maintain a nominal spin about the $\underrightarrow{b}^{3}$ axis, this choice allows for ease of linearization about any nominal angular velocity about the $\underrightarrow{b}^{3}$ axis, and positions the kinematic singularity at $180\text{\textdegree}$ from the nominal inertial pointing attitude.

### v2

@@ Abstract
This paper presents a novel disturbance-torque-estimation-augmented model predictive control (MPC) framework to perform momentum management on NASA’s Solar Cruiser solar sail mission. Solar Cruiser represents a critical step in the advancement of large-scale solar sail technology and includes the innovative use of an active mass translator (AMT) and reflectivity control devices (RCDs) as momentum management actuators. The coupled nature of these actuators has proven challenging in the development of a robust momentum management controller. Recent literature has explored the use of MPC for solar sail momentum management with promising results, although exact knowledge of the disturbance torques acting on the solar sail was required. This paper amends this issue through the use of a Kalman filter to provide real-time estimation of unmodeled disturbance torques. Furthermore, the dynamics model used in this paper incorporates key fidelity enhancements compared to prior work, including Solar Cruiser’s four-reaction-wheel assembly and the offset between its center of mass and center of pressure. More realistic operation scenarios involving the tracking of large angle slew maneuvers under attitude-dependent solar radiation force and torque are also performed to further validate the proposed method compared to prior work. Simulation results demonstrate that the proposed policy successfully manages angular momentum growth under slew maneuvers that exceed the operational envelope of the current state-of-the-art method. The inclusion of the disturbance torque estimate is shown to greatly improve the reliability and performance of the proposed MPC approach. This work establishes a new benchmark for Solar Cruiser’s momentum management capabilities and paves the way for MPC-based momentum management of other solar sails making use of an AMT and/or RCDs.
@@ 1 Introduction
Prior work by Shen and Caverly (2026) developed MPC-based momentum management strategies tailored for solar sails equipped with an AMT and RCDs. They developed a dynamics model that captured key features of the AMT movement, including the resultant time-varying changes in the spacecraft’s CM and moment of inertia matrix. The MPC policy developed in the work of Shen and Caverly (2026) leveraged its optimization capabilities to handle the actuator constraints and requirements, including the enforcement of on-off RCD actuation and AMT motion rate limits, all while incorporating tuning parameters designed to adjust the trade-off between system performance and control effort. For real-time onboard implementation, the MPC formulation was posed as a quadratic program (QP), which guarantees fast and robust convergence suitable for the short processing cycles required by the flight computer. The work of Shen and Caverly (2025) further examined the critical balance between model fidelity and computational cost when implementing the MPC policy developed by Shen and Caverly (2026) to determine feasibility for real-time implementation. However, these MPC implementations both make unrealistic assumptions that the SRP force and torque acting on the solar sail remain constant and that exact knowledge of the disturbance torque acting on the solar sail is available for use within MPC’s prediction model. These are significant assumptions that limit the practical implementation of the MPC policy proposed by Shen and Caverly (2026). The SRP forces and torques acting on the solar sail are attitude-dependent and it is virtually impossible to accurately predict disturbance torques from analytical models due to the unpredictable shape deformation of the solar sail and temporal changes in the sail’s optical properties (Wang et al., 2025; Gauvain and Tyler, 2023). The use of an inaccurate torque model within the MPC framework significantly degrades momentum management performance, negating the purported benefits of the MPC momentum management policy. Another practical limitation of the work of Shen and Caverly (2026, 2025) is that their implementations assume that the solar sail is equipped with three RWs aligned with the principal axes of its body-fixed frame. Many spacecraft, including Solar Cruiser, have a 4-RW assembly for redundancy and increased performance, which precludes the use of the methods developed by Shen and Caverly (2026, 2025). Furthermore, the implementation in the work of Shen and Caverly (2026) assumed that the RWs remained in the same plane as the solar sail’s CP, which is not representative of Solar Cruiser’s geometry. A final limitation to note in the work of Shen and Caverly (2026) is that it is only capable of regulating the solar sail to a fixed attitude (i.e., an attitude hold). This does not meet the practical needs of a solar sail mission, which may require performing attitude slews.
@@
To overcome the limitations of prior work, this paper presents a novel MPC-based momentum management policy that incorporates disturbance torque estimation, a 4-RW assembly tailored for the Solar Cruiser mission, an attitude-dependent SRP force and torque, and the ability to track attitude slews. A Kalman filter framework is used to estimate the unmodeled disturbance torques and system model errors in real time, thus enhancing the predictive capability of the MPC. Similar Kalman-filtering approaches have been used in the literature to estimate unknown parameters or terms within a system model (Zenere and Zorzi, 2018; Woodbury and Junkins, 2010; Hayes and Caverly, 2025; Ahmed et al., 2024). For example, Hayes and Caverly (2025) used a Kalman filter to estimate the unknown atmospheric density of a satellite during an orbital reentry, while Ahmed et al. (2024) estimated the unknown wind acting on a small uncrewed air vehicle. Solar Cruiser’s 4-RW assembly is accounted for within the proposed MPC implementation through the use of the commonly-used pseudo-inverse RW allocation approach (Leve et al., 2015; Markley and Crassidis, 2014). This provides the MPC prediction model with accurate knowledge of the dynamics of each individual RW, allowing for their operation to be constrained within their saturation limits.
@@
This paper presents four key contributions relative to the state-of-the-art in solar sail momentum management, including prior work on MPC-based methods by Shen and Caverly (2026, 2025). The first contribution is a robust MPC momentum management formulation that uses a Kalman filter to estimate unknown disturbance torques acting on the solar sail. To the best of the knowledge of the authors, this is the first realistically-implementable momentum management policy for a solar sail equipped with an AMT and RCD that outperforms the method of Tyler et al. (2023). The second contribution is the incorporation of a 4-RW assembly within an MPC-based momentum management policy. To the best of the knowledge of the authors, this is the first MPC-based momentum management policy to consider a realistic 4-RW assembly. The third contribution is an assessment of the proposed momentum management policy in a realistic simulation of Solar Cruiser’s dynamics. Specifically, its CM is located a distance from the sail plane, non-ideal and attitude-dependent SRP forces and torques are considered, and the magnitude of the roll torque generated by the RCDs is modeled as attitude-dependent. All of these effects are meaningful when considering Solar Cruiser’s dynamics, as they result in substantial coupling between the system’s dynamics and actuation, yet they were not considered in the prior work of Shen and Caverly (2026). The fourth contribution is the augmentation of the proposed MPC policy to track large angle slew maneuvers, which is an important operational requirement of a typical solar sail mission. This improvement expands well-beyond the attitude hold capabilities formulated and demonstrated in the prior work of Shen and Caverly (2026).
@@
Details of the nonlinear system dynamics of Solar Cruiser, the 4-RW control allocation algorithm, and the momentum management actuators are presented in Section 2. This section also provides the linearized dynamics model used in the Kalman filter and MPC frameworks. The Kalman filter formulation is presented in Section 3, providing details of how the unmeasurable disturbance torques are estimated. The MPC formulation is presented in Section 4, detailing the implementation of estimation-prediction framework and the incorporation of the 4-RW assembly into the MPC prediction model. Numerical simulation results are presented in Section 5, validating the performance of the proposed estimation-augmented MPC with comparisons to Solar Cruiser’s state-of-the-art momentum management method (Tyler et al., 2023). Results in this section are also presented that demonstrate the effect of actuation thresholds within the proposed momentum management policy on actuation efficiency and observability of the roll-axis disturbance torque. A description of the reference slew maneuver used in this work is included in the Appendix.
@@ 2 Attitude Dynamics and Control Actuation
NASA’s Solar Cruiser uses AMT and RCDs as its momentum management actuators (Inness et al., 2023). The AMT changes the relative alignment of the CM and CP such that the SRP force acting on the CP results in a corresponding torque about the CM. Moving the AMT and appropriately placing the CM/CP offset results in a controllable moment that unloads the accumulated RW angular momentum. However, the moment of inertia matrix changes when the AMT moves and the mass distribution of the sailcraft changes. The dynamics are thus coupled with the AMT translation. This section presents the dynamics and control of the sailcraft, starting with important notation and proceeding with its attitude dynamics, the RW attitude control law, details regarding the momentum management control actuation, and the linearized attitude dynamics model used for the Kalman filter and MPC frameworks.
@@ 2.1 Notation
The direction cosine matrix (DCM) $\mbf{C}_{ba}$ describes the attitude of reference frame $\mathcal{F}_{b}$ relative to reference frame $\mathcal{F}_{a}$. While different attitude parameterizations can be used to describe a DCM, Euler-angle sequences are used in this work due to their ease of physical interpretation and the lack of any large-angle maneuvers that would potentially result in a kinematic singularity. The DCM can be used to express a physical vector in different reference frames. For example, $\mbf{v}_{b}=\mbf{C}_{ba}\mbf{v}_{a}$.
@@ 2.2 Solar Cruiser Attitude Dynamics
Following the approach of Shen and Caverly (2026) and as illustrated in Fig. 1, Solar Cruiser’s sail is modeled as a thin flat plate (denoted as $\mathcal{S}$) and a rigid rectangular bus (denoted as $\mathcal{P}$). The nonlinear rigid-body attitude dynamics of a solar sail incorporating the AMT translation as a control input were derived by Shen and Caverly (2026), including the time-dependent moment of inertia matrix corresponding to the AMT position. It is assumed that the each RW spin axis aligns with the position of the RW relative to the CM of the bus (point $p$). Although this assumption is not necessarily required to implement the proposed momentum management method, it removes additional coupling terms in the dynamics that can often be calibrated for upon spacecraft commissioning. As shown in Fig. 1, the sailcraft body frame $\mathcal{F}_{b}$ is defined with $\underrightarrow{b}^{3}$ pointing through the normal (roll) axis of the sail, and $\underrightarrow{b}^{2}$, $\underrightarrow{b}^{1}$ pointing within the plane of the sail and representing the pitch and yaw axes, respectively. The moment of inertia matrix relative to the sailcraft’s CM (point $c$) is defined as
@@
Let $\mathcal{F}_{a}$, defined by basis vectors $\underrightarrow{a}^{1}$, $\underrightarrow{a}^{2}$, $\underrightarrow{a}^{3}$, be an inertial reference frame. The solar sail attitude dynamics with mass translation are defined as (Shen and Caverly, 2026)
@@
where ${\boldsymbol{\omega}}_{b}^{ba}$ is the angular velocity of $\mathcal{F}_{b}$ relative to $\mathcal{F}_{a}$ resolved in $\mathcal{F}_{b}$, $\mbf{h}_{b}^{\text{RW}}$ is the collective angular momentum of the RWs relative to the CM of the bus (point $p$) with respect to inertial reference frame $\mathcal{F}_{a}$ resolved in $\mathcal{F}_{b}$. The term ${\boldsymbol{\tau}}_{b}^{\mathcal{B}c}={\boldsymbol{\tau}}_{b}^{\text{AMT}}+{\boldsymbol{\tau}}_{b}^{\text{RCD}}+{\boldsymbol{\tau}}_{b}^{\text{dist}}$ is the torque acting on the sailcraft relative to its CM, consisting of the AMT-SRP-induced torque ${\boldsymbol{\tau}}_{b}^{\text{AMT}}=\frac{m_{s}}{m_{p}+m_{s}}\mbf{r}_{b}^{ps^{\times}}\mbf{f}_{b}^{\text{SRP}}$, the RCD torque ${\boldsymbol{\tau}}_{b}^{\text{RCD}}$, and the disturbance torque ${\boldsymbol{\tau}}_{b}^{\text{dist}}$. The term $\mbf{f}_{b}^{\text{SRP}}$ denotes the solar radiation pressure induced force acting on the sail’s CP. The RCD torque is generated by activating one of the set of four RCDs (either brown or orange) visualized in Fig. 1. Actuating the orange RCDs generates a positive roll torque about $\underrightarrow{b}^{3}$ axis, while actuating the brown RCDs generates a negative torque about this same axis. The time-dependent argument (t) for the variables $\mbf{J}_{b}^{\mathcal{B}c}(t)$, $\mbf{r}^{ps}_{b}(t)$, ${\boldsymbol{\omega}}_{b}^{ba}(t)$, $\mbf{h}_{b}^{\text{RW}}(t)$, ${\boldsymbol{\tau}}_{b}^{\mathcal{B}c}(t)$ is omitted for brevity in Eq. (1).
@@
A 3-2-1 Euler angle sequence is used to describe the rotation between $\mathcal{F}_{a}$ and $\mathcal{F}_{b}$, so that $\mbf{C}_{ba}=\mbf{C}_{1}(\theta_{1})\mbf{C}_{2}(\theta_{2})\mbf{C}_{3}(\theta_{3})$ is the DCM describing the orientation of $\mathcal{F}_{b}$ relative to $\mathcal{F}_{a}$, where $\mbf{C}_{i}(\cdot)$ is the DCM representing a rotation about the $i$-th principal axis. The mapping ${\boldsymbol{\omega}}^{ba}_{b}=\mbf{S}({\boldsymbol{\theta}})\dot{{\boldsymbol{\theta}}}$ relates Euler angle rates to angular velocity, where the column matrix ${\boldsymbol{\theta}}=\Big[\theta_{1}\,\,\,\theta_{2}\,\,\,\theta_{3}\Big]^{\mathsf{T}}$ is the set of Euler angles and $\mbf{S}({\boldsymbol{\theta}})$ is the mapping matrix. It is worth noting that $\mbf{S}({\boldsymbol{\theta}})$ depends only on $\theta_{1}$ and $\theta_{2}$ due to the selected 3-2-1 Euler angle sequence. Given that a solar sail is designed to keep the Sun within its field of view and maintain a nominal spin about the $\underrightarrow{b}^{3}$ axis, this choice allows for ease of linearization about any nominal angular velocity about the $\underrightarrow{b}^{3}$ axis, and positions the kinematic singularity at $180\text{\textdegree}$ from the nominal inertial pointing attitude.
Generally, the forces and torques induced by SRP vary with the attitude of the solar sail, the shape of the sail membrane, and its optical properties. Given the slow nature of solar sail maneuvers, it is assumed that vibrations in the structure are minimal and both the sail shape and material properties remain static, which implies that $\mbf{f}_{b}^{\text{SRP}}$ and ${\boldsymbol{\tau}}_{b}^{\text{dist}}$ depend solely on the attitude of the sailcraft. The optical properties of the Solar Cruiser membrane (Heaton and Artusio-Glimpse, 2015; Heaton et al., 2017) and its expected deformations (Gauvain and Tyler, 2023) are incorporated into the static membrane shape model developed by Bunker and Caverly (2026). This model is then employed to compute $\mbf{f}_{b}^{\text{SRP}}$ and ${\boldsymbol{\tau}}_{b}^{\text{dist}}$ based on the solar sail’s attitude $\boldsymbol{\theta}$} in this work.